OpenStudy (anonymous):

5a^2(3a^4+3b+2) ^ = exponent. Please help! :D

OpenStudy (raden):

use the distributive property : a(b+c+d) = ab + ac + ad

OpenStudy (raden):

and exponent property : a^m * a^n = a^(m+n)